# Fine-grained Stateful Knowledge Exploration: Effective and Efficient Graph Retrieval with Large Language Models

### v3

@@ Abstract
Large Language Models (LLMs) have shown impressive capabilities, yet updating their knowledge remains a significant challenge, often leading to outdated or inaccurate responses. A proposed solution is the integration of external knowledge bases, such as knowledge graphs, with LLMs. Most existing methods use a paradigm that treats the question as the objective, with relevant knowledge being incrementally retrieved from the knowledge graph. However, this strategy frequently experiences an mismatch in the granularity of knowledge between the target question and the entities and relations being retrieved. As a result, the information in the question cannot precisely correspond to the retrieved knowledge. This may cause redundant exploration or omission of vital knowledge, thereby leading to enhanced computational consumption and reduced retrieval accuracy. In this paper, we propose a novel paradigm of fine-grained stateful knowledge exploration, which addresses the ‘information granularity mismatch’ issue. We extract fine-grained information from questions and explore the semantic mapping between this information and the knowledge in graph. By dynamically updating the mapping records, we avoid redundant exploration and ensure no pertinent information is overlooked, thereby reducing computational overhead and improving the accuracy of knowledge exploration. The use of fine-grained information also eliminates the need for a priori knowledge, a common requirement in existing methods. Experiments on multiple datasets revealed that our paradigm surpasses current advanced methods in knowledge retrieval while significantly reducing the average number of LLM invocations. The code for this paper can be found at https://github.com/nnnoidea/stateful-KGQA.
@@ 1 Introduction
Large language models (LLMs) have demonstrated considerable efficacy in a range of natural language processing tasks, as evidenced by recent studies Achiam et al. (2023); Cheng et al. (2023); Meta (2024); Zhao et al. (2024). Nevertheless, there are certain limitations to be considered when employing LLMs in question-answering tasks. The knowledge of LLMs remains static, fixed at the point of their training, which presents a challenge in incorporating timely updates. Moreover, in response to queries with which they are unfamiliar, LLMs may occasionally provide incorrect answers accompanied by an overconfident tone Luo et al. (2023), a phenomenon that is often referred to as hallucination Bang et al. (2023); Ji et al. (2023). In light of these challenges, an effective solution is the integration of external knowledge, specifically knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
@@
A number of studies have been conducted with the objective of combining external knowledge with LLMs. One straightforward approach is to prompt LLMs to directly generate the corresponding knowledge base operation statements based on questions. This method has been demonstrated to be effective in various studies, as referenced in Hu et al. (2023); Wang et al. (2023). Additionally, chain-of-thought Wei et al. (2022b) is employed to assist LLMs in decomposing input questions into multiple sub-questions Li et al. (2023a); Zhao et al. (2023). However, these methods do not take into account the awareness of the nature of external knowledge sources on the part of LLMs. Consequently, the instructions or sub-questions generated by LLMs may not be aligned with the external knowledge. To avoid the mismatch issue due to unawareness of the knowledge base, current methods no longer try to get the entire knowledge path in a single step. Instead, they acquire it incrementally through iteration. These methods furnish the LLM with subgraph information from the knowledge base, thereby averting the occurrence of mismatch during each iteration, such as Sun et al. (2024).
However, the existing methods only avoid the mismatch at the character level without noticing the mismatch at the information granularity. The question, as a sentence containing multiple pieces of information, has an information granularity that is significantly larger than the knowledge granularity of entities and relations in the knowledge graph. This leads to two problems. First, the information in question does not correspond precisely to the retrieved knowledge, so existing methods have to engage in stateless iterative exploration. In stateless iterative exploration, the same sentence is matched with subgraphs containing different entities and relations in different rounds. It is important to note that some information in the sentence may be matched repeatedly in different rounds, while other information may be ignored from the beginning to the end. An example is shown in the upper half of the Fig. 1. On the one hand, repeated matching increases the number of exploration rounds, resulting in higher computational consumption, especially for large language models. On the other hand, if some information is ignored, the retrieved knowledge will be incomplete and the question-answering effect will be compromised. Second, it is evident that current methodologies require the artificial provision of fine-grained topic-entity as a starting point for knowledge exploration paths Sun et al. (2024); Chen et al. (2024). This will undoubtedly result in additional manual effort in practice. Some approaches are vaguely aware of the information granularity mismatch problem and split the query into multiple clauses Chen et al. (2024), but do not fundamentally address the mismatch in information granularity.
To solve the above problems caused by the mismatch of information granularity, in this paper we propose a new paradigm of fine-graind stateful knowledge exploration, as shown in the lower part of Fig. 1. The new paradigm extracts fine-grained, sufficient information from the question to capture most of its semantics. It also transforms the goal of knowledge exploration from coarse-grained to multiple independent fine-grained information. We call this fine-grained information clues. When performing knowledge exploration, the knowledge corresponding to a clue is explored iteratively from the knowledge graph. We call this correspondence between the clues and the knowledge in the graph a mapping. In this stateful knowledge exploration process, keeping track of the clues and their mappings ensures that different rounds of exploration do not explore duplicate information, and also avoids ignoring some information. An example of this is shown in Figure 1. In this new paradigm, using clues as the starting point of knowledge exploration eliminates the need for human-provided topic entities, a common requirement in existing methods.
@@
  • We point out the issue of information granularity mismatch in the existing mainstream paradigms, which forces current methods to perform stateless iterative exploration. This results in increased computational consumption in the knowledge exploration process, incomplete retrieval of knowledge, and additional requirements for prior knowledge.
  • We propose a new paradigm for stateful knowledge exploration based on fine-grained information, which resolves the issue of information granularity mismatch. The new paradigm eliminates the need for prior knowledge. Stateful iterative exploration significantly reduces the computational consumption in the knowledge retrieval process and improves the effectiveness of knowledge retrieval.
  • We conducted experiments on multiple datasets. The results show that our proposed new paradigm performs excellently on all datasets. The new paradigm not only outperforms existing methods in knowledge graph question-answering but also significantly reduces the number of LLM calls, sometimes to only a fraction of that required by existing methods.
@@ 2.1 Document Retrieval Enhanced Models
2.2 Knowledge base Enhanced Models
There are diverse ways to enhance models based on knowledge, aiming to address issues arising from complex reasoning, hallucination, and poor interpretability. Cao et al Cao et al. (2021) improved implicit event causality extraction by leveraging structured knowledge bases. Zhao et al Zhao et al. (2023) recently proposed a Verification and Editing (VE) framework. This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systems Wei et al. (2022a). Researchers have also introduced the Chain-of-Knowledge (CoK) Li et al. (2023b), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources. This ensures that inaccuracies in the foundational principles do not propagate to subsequent steps. Think-on-Graph(ToG) Sun et al. (2024) is a plug-and-play framework that seamlessly applies to various LLMs and KGs. ToG also enhances the reasoning capabilities, enabling smaller LLMs to compete with larger LLMs.
@@ 3.1 Problem definition
Knowledge Graph Question-Answering is the task of retrieving relevant knowledge from a knowledge graph to answer given questions. In our proposed new paradigm, we first obtain a set of clues $clue_{all}=\{\text{clue}_{1},\text{clue}_{2},\dots\}$ from the question q. For each clue in the $clue_{all}$, we explore its corresponding relations $r$ and entities $e$ in the knowledge graph, forming a knowledge path $P$ = $\{(e^{0},r^{1},e^{1}),(e^{1},r^{2},e^{2}),\dots,(e^{i-1},r^{i},e^{i})\}$. This path then serves as a base for the large language model to answer the question.
3.2 Fine-grained clues extraction
To achieve stateful knowledge exploration, we need to obtain fine-grained information, which we refer to as clues. Clues should not only characterize the information of the question, but also match the granularity of information, or information content, with the entities and relations in the knowledge graph. Intuitively, we posit that the majority of information (not all information) within a question can be characterized by keywords originating from the question. Simultaneously, the granularity of keywords is comparable to that of entities in the knowledge graph. The use of these keywords as clues for knowledge retrieval helps in addressing the given question. For example, in the question ‘What is the population of the capital of China?’, the key information required is ‘population’, while ‘China’ and ‘capital’ serve to narrow the context. (China, capital, population) are the keywords in the given question. However, it is important to note that different knowledge graphs have different ontology designs. The same clue may correspond to an entity in graph A while being related to a relation in graph B. To heighten the universality of our approach, we refrain from making explicit distinctions between clues that map to entities or relations during the clue extraction phase. Instead, we simply prompt the LLM to extract the entire set of clues.
@@ 3.3 Starting point
After obtaining the fine-grained clues, we first determine the starting point of the exploration. Unlike existing methods that require human-provided topic entities as starting points, the clues we extract can serve directly as starting points. It was mentioned above that some clues may map to relations, but information extracted from a semantically complete sentence necessarily has one or more mappings to entities, such as ‘China’ in the example above. Sentences without entities cannot effectively express semantics. This is not only consistent with our knowledge, but also verified in our experiments. We retrieve entities that match the clue characters in the knowledge graph and use any successful matching entity as a starting point. Then, we will gradually explore the entire knowledge path based on this starting point.
3.4 Stateful knowledge path exploration
Through iterative exploration, we explore the entities or relations mapping to each clue, forming a knowledge path. Each round of mapping exploration consists of two steps: relation-based mapping exploration and entity-based mapping exploration. For better understanding, we describe the process using the $n$-th iteration as an example. In the previous $n-1$ iterations, we have obtained $n-1$ triples $(e^{i-1},r^{i},e^{i})$, where $e^{i-1}$ is the tail entity of the $(i-1)$-th triple and the head entity of the $i$-th triple, and $r^{i}$ is the relation of the $i$-th triple. The starting point $e^{0}$ is the head entity of the first triple. The set of clues already explored is $C$, and the remaining set of clues $C^{\prime}=clue_{all}-C$.
@@ 3.4 Stateful knowledge path exploration
The first step in each round of mapping exploration is relational mapping exploration. In the $n$-th iteration, we first obtain the relations $R=\{r_{1},r_{2},r_{3},r_{4},...\}$ associated with the entity $e_{n-1}$ from the graph, where $e^{n-1}$ is the tail entity of the triple obtained in the $(n-1)$-th exploration. We provide $R$ and $C^{\prime}$ to LLM and ask LLM to determine if there exists a $clue_{i}\in C^{\prime}$ that maps to any $r$. The mapping result is denoted as:
@@
Three different results can occur:
@@
[…]
3. e-mapping is empty and then terminate the exploration.
3.5 Ending exploration and answering questions
Every step in the exploration process might result in multiple outcomes, thus generating multiple branches. To prevent an explosion in the number of branches, existing methods opt to select the top-N branches as final results. However, this approach could inevitably lead to omissions, particularly when the number of correct results exceeds N. In our method, we do not filter branches based on the quantity, but instead, on whether all clues can be mapped. Therefore, during the exploration process, any branch unable to find the next clue mapping will be terminated. If a branch that finds all clue mappings emerges, then the exploration of all other branches concludes simultaneously. At the end of the exploration, if there exists a knowledge path where all clue mappings have been found, it is offered to the large language model to answer the question. If no such path exists, then it is assumed that the exploration failed, and the question is directly posed to the large language model for an answer.
Table 1: Results for MOOC Q&A.

### v4

@@ Abstract
Large Language Models (LLMs) have shown impressive capabilities, yet updating their knowledge remains a significant challenge, often leading to outdated or inaccurate responses. A proposed solution is the integration of external knowledge bases, such as knowledge graphs, with LLMs. Most existing methods use a paradigm that treats the whole question as the objective, with relevant knowledge being incrementally retrieved from the knowledge graph. However, this paradigm often leads to a granularity mismatch between the target question and the retrieved entities and relations. As a result, the information in the question cannot precisely correspond to the retrieved knowledge. This may cause redundant exploration or omission of vital knowledge, thereby leading to enhanced computational consumption and reduced retrieval accuracy. To address the limitations of coarse-grained knowledge exploration, we propose FiSKE, a novel paradigm for Fine-grained Stateful Knowledge Exploration. FiSKE first decomposes questions into fine-grained clues, then employs an adaptive mapping strategy during knowledge exploration process to resolve ambiguity in clue-to-graph mappings. This strategy dynamically infers contextual correspondences while maintaining a stateful record of the mappings. A clue-driven termination mechanism ensures rigorous augmentation—leveraging fully mapped paths for LLMs while reverting to chain-of-thought reasoning when necessary. Our approach balances precision and efficiency. Experiments on multiple datasets revealed that our paradigm surpasses current advanced methods in knowledge retrieval while significantly reducing the average number of LLM invocations. The code for this paper can be found at https://github.com/nnnoidea/stateful-KGQA.
keywords:
[…]
{highlights}
First to systematically diagnose the granularity mismatch problem in mainstream KG-LLM paradigms, revealing its threefold impact: computational redundancy, incomplete retrieval, and unnecessary prior knowledge dependence.
Propose FiSKE, the first stateful fine-grained KG exploration framework, enabling atomic question decomposition, cross-iteration state tracking, and dynamic switching between KG augmentation and chain-of-thought reasoning.
Resolve clue-to-KG mapping ambiguity via a novel adaptive strategy that dynamically infers contextual correspondences (entity/relation/both), eliminating redundant exploration while ensuring retrieval precision.
@@ 1 Introduction
Large language models (LLMs) have demonstrated considerable efficacy in a range of natural language processing tasks, as evidenced by recent studies Achiam et al. (2023); Cheng et al. (2023); Meta (2024); Zhao et al. (2024); Li et al. (2024). Nevertheless, there are certain limitations to be considered when employing LLMs in question-answering tasks. The knowledge of LLMs remains static, fixed at the point of their training, which presents a challenge in incorporating timely updates. Moreover, in response to queries with which they are unfamiliar, LLMs may occasionally provide incorrect answers accompanied by an overconfident tone Luo et al. (2023), a phenomenon that is often referred to as hallucination Bang et al. (2023); Ji et al. (2023). In light of these challenges, an effective solution is the integration of external knowledge Nie et al. (2020), specifically knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
Figure 1: Compared to existing methods, our approach exhibits two key distinctions: (1) no manual provision of topic entities is required, and (2) it eliminates both information oversight and redundant exploration.
A number of studies have been conducted with the objective of combining external knowledge with LLMs. One straightforward approach is to prompt LLMs to directly generate the corresponding knowledge base operation statements based on questions. This method has been demonstrated to be effective in various studies, as referenced in Hu et al. (2023); Wang et al. (2023b). Additionally, chain-of-thought Wei et al. (2022) is employed to assist LLMs in decomposing input questions into multiple sub-questions Li et al. (2023b); Zhao et al. (2023). However, these methods do not take into account the awareness of the nature of external knowledge sources on the part of LLMs. Consequently, the instructions or sub-questions generated by LLMs may not be aligned with the external knowledge. To avoid the mismatch issue due to unawareness of the knowledge base, ToG Sun et al. (2024) leverages the semantic capabilities of LLMs to retrieve relevant knowledge from the knowledge graph.
However, existing methods primarily focus on addressing mismatches at the character level, while largely overlooking the more critical issue of mismatches at the information granularity level. A question, as a sentence comprising multiple pieces of information, inherently exhibits a coarser information granularity compared to the fine-grained entities and relations in the knowledge graph. As shown in the upper half of Fig. 1, this discrepancy gives rise to the following two major challenges.
First, existing methods can only verify the relevance of retrieved knowledge (entities/relations) to the question, but not their redundancy or completeness. Consequently, current approaches are constrained to stateless iterative exploration (i.e., processing each iteration independently without retaining cross-iteration information), wherein the identical sentence must be repeatedly reprocessed across iterations, each time realigning with newly retrieved subgraphs containing different entity-relation combinations. This stateless nature inherently leads to two significant inefficiencies: redundant exploration and information omission. Specifically, certain pieces of information in the sentence may be redundantly explored across different rounds, while others may remain entirely overlooked throughout the process. Such limitations not only compromise the accuracy of knowledge retrieval but also result in suboptimal computational efficiency. Notably, while some approaches attempt to mitigate granularity mismatch through query splitting Chen et al. (2024), these patchwork solutions fail to address the root cause: the persistent disparity between question-level abstraction and KG-level granularity. Second, current methodologies require manual provision of fine-grained topic entities as exploration starting points Sun et al. (2024); Chen et al. (2024), inevitably demanding additional human effort in practical applications. These dual challenges highlight the need for a state-aware approach capable of dynamically bridging question contexts with KG knowledge.
To solve the above problems caused by the mismatch of information granularity, in this paper we propose a novel paradigm of Fine-grained Stateful Knowledge Exploration, which we refer to as FiSKE, as illustrated in the lower part of Fig. 1. This new paradigm extracts fine-grained, sufficient information from the question to capture most of its semantics. It also shifts the goal of knowledge exploration from coarse-grained to multiple independent fine-grained information units, which we refer to as clues. During knowledge exploration, the knowledge corresponding to each clue is iteratively retrieved from the knowledge graph. We define this correspondence between the clues and the knowledge in the graph as a mapping. By maintaining a stateful record of the clues and their mappings across iterations, FiSKE ensures that different rounds of exploration avoid redundant information retrieval while preventing the oversight of critical details. With its clue-driven termination mechanism, FiSKE employs only fully mapped knowledge paths for LLM augmentation. If no valid paths are found, it automatically falls back to chain-of-thought reasoning. This hybrid approach ensures both rigor and efficiency in knowledge utilization. A simple path exploration example is shown in the Fig. 2, demonstrating the impact of our proposed stateful method on avoiding redundant exploration.
However, this fine-grained paradigm introduces a unique challenge in mapping resolution: the extracted clues may correspond to different structural elements in the knowledge graph. Specifically, while some clues distinctly map to either entities or relationships, others may simultaneously satisfy the conditions for both. This ambiguity arises precisely because our fine-grained method is designed to evaluate the redundancy or completeness of retrieved knowledge, a capability absent in existing approaches. Current methods rely on coarse question-entity mapping, which lacks mechanisms to assess knowledge redundancy/completeness, let alone resolve such ambiguous mappings. To address this, we introduce an adaptive mapping strategy that dynamically resolves whether a clue corresponds to an entity, a relationship, or both, based on contextual evidence from the knowledge graph.
@@
  • We identify and highlight the critical issue of information granularity mismatch in existing mainstream paradigms. This mismatch forces current methods to rely on stateless iterative exploration, leading to increased computational overhead during knowledge exploration, incomplete retrieval of knowledge, and unnecessary reliance on prior knowledge. Our analysis provides a deeper understanding of the limitations inherent in these approaches.
  • We propose FiSKE, a novel paradigm that transforms knowledge exploration into a stateful, fine-grained process. By decomposing questions into atomic clues and maintaining cross-iteration exploration states, FiSKE eliminates redundancy and omissions inherent in stateless approaches. The paradigm further integrates a clue-driven termination mechanism to dynamically balance KG augmentation and chain-of-thought reasoning.
  • We identify a key challenge in fine-grained knowledge exploration: extracted clues may have ambiguous mappings to knowledge graph elements, potentially corresponding to entities, relations, or both. To resolve this mapping ambiguity, we propose an adaptive mapping strategy that dynamically determines the most appropriate mapping for each clue based on contextual evidence from the knowledge graph.
  • We conduct extensive experiments on multiple datasets to validate the effectiveness of our proposed paradigm. The results demonstrate that our method achieves superior performance across all datasets, outperforming existing methods in knowledge graph question-answering tasks. Additionally, our paradigm drastically reduces the number of large language model calls, often reducing them to a fraction of those required by current approaches, thereby improving computational efficiency.
@@ 2.1 Document Retrieval Enhanced Models
2.2 Knowledge Graph Enhanced Models
There are diverse ways to enhance models based on knowledge, aiming to address issues arising from complex reasoning, hallucination, and poor interpretability. Cao et al Cao et al. (2021) improved implicit event causality extraction by leveraging structured knowledge bases. Zhao et al Zhao et al. (2023) recently proposed a Verification and Editing (VE) framework. This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systems Wei et al. (2022). Researchers have also introduced the Chain-of-Knowledge (CoK) Li et al. (2023c), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources. This ensures that inaccuracies in the foundational principles do not propagate to subsequent steps. Think-on-Graph(ToG) Sun et al. (2024) is a plug-and-play framework that seamlessly applies to various LLMs and KGs. ToG also enhances the reasoning capabilities, enabling smaller LLMs to compete with larger LLMs.
@@ 3.1 Problem Definition
Knowledge Graph Question-Answering is the task of retrieving relevant knowledge from a knowledge graph to answer given questions. In our proposed new paradigm, we first obtain a set of clues $clue_{all}=\{\text{clue}_{1},\text{clue}_{2},\dots\}$ from the question q. For each clue in the $clue_{all}$, we explore its corresponding relations $r$ or entities $e$ in the knowledge graph, forming a knowledge path $P$ = $\{(e^{0},r^{1},e^{1}),\dots,(e^{i-1},r^{i},e^{i})\}$. This path then serves as a base for the large language model to answer the question.
3.2 Extraction of Fine-grained Clues
To enable stateful knowledge exploration, it is essential to extract fine-grained information, which we refer to as clues. We posit that the majority of information within a question can be effectively characterized by its keywords, whose granularity is comparable to that of entities in the knowledge graph. Using these keywords as clues facilitates knowledge retrieval and enables the resolution of the given question.
For example, in the question “What is the population of the capital of China?", the key information required is “population", while “China" and “capital" provide contextual constraints. Thus, the keywords extracted from this question are (China, capital, population), which serve as the clues for knowledge exploration.
Clues may correspond to entities, relations, or their combinations. We leverage the capabilities of the LLM to systematically extract the complete set of clues, resulting in an ontology-agnostic approach. Example prompt can be seen in Appendix A.1. The exact mappings are resolved dynamically during path exploration through our adaptive mapping strategy.
@@ 3.3 Starting Point
After extracting the fine-grained clues, our approach uses these clues to identify exploration starting points in the knowledge graph. Unlike existing methods that require human-provided topic entities, we automatically determine starting points through exact lexical matching (e.g., using a basic database operation like SELECT or MATCH). Since well-formed questions typically contain at least one identifiable entity reference, this matching process reliably anchors the exploration. Each successfully matched entity serves as a starting point, from which we iteratively explore the complete knowledge path.
3.4 Adaptive and Stateful Knowledge Exploration
Through iterative exploration, we map clues to corresponding entities or relations, constructing a knowledge path $P$. We maintain a stateful record of these mappings to dynamically adjust candidate clues in each round. An adaptive mapping strategy determines whether current-step triples should map to single or multiple clues.
The strategy comprises two steps: relation-based and entity-based mapping exploration. Using the $n$-th iteration as an example, previous iterations yield $n-1$ triples forming path $P=\{(e^{0},r^{1},e^{1}),\dots,(e^{n-2},r^{n-1},e^{n-1})\}$, where $e^{0}$ is the initial head entity. The explored clue set $C$ prevents redundant processing, with candidate clues $C^{\prime}=clue_{all}-C$.
@@ 3.4 Adaptive and Stateful Knowledge Exploration
In the $n$-th iteration, we first retrieve relations $R=\{r_{1},\dots,r_{m}\}$ for entity $e^{n-1}$, then prompt LLM to map $C^{\prime}$ to $R$, yielding:
@@
Three outcomes occur:
Case 1: Single-clue dual mapping $(\text{clue}_{1},r_{i})$ and $(\text{clue}_{1},e_{j})$:
  • Update $C\leftarrow C\cup\{clue_{1}\}$
  • Extend $P\leftarrow P\cup\{(e^{n-1},r_{i},e_{j})\}$
Case 2: Dual-clue mapping $(\text{clue}_{1},r_{i})$ and $(\text{clue}_{2},e_{j})$:
  • Update $C\leftarrow C\cup\{clue_{1},clue_{2}\}$
  • Extend $P$ as in Case 1
@@
The strategy iteratively constructs $P$ through LLM mapping with stateful tracking of $C$ and $C^{\prime}$. Example prompts appear in Appendix A.1. The pseudocode can be found in Appendix A.2.
3.5 Clue-Driven Termination and Branch-Adaptive Answering
During the exploration process, each step may generate multiple outcomes, resulting in multiple branches. To prevent an exponential increase in the number of branches, existing methods typically select the top-N branches as final results. However, this approach inevitably risks omissions, especially when the number of correct results exceeds N. In our method, we do not filter branches based on quantity but instead rely on whether all clues can be successfully mapped. Specifically, for any branch, if it fails during the mapping exploration phase, we deem it incapable of fully resolving all clues and thus discard it from further consideration. Conversely, if a branch successfully maps all clues, the exploration is considered complete. Upon finishing the current round of exploration across all branches, the process terminates, and no further exploration is conducted.
If a branch successfully completes all mappings, we provide its knowledge path $P$ as external knowledge to the LLM, instructing the model to answer questions based on this contextual information. However, if all branches fail to achieve complete mapping, we infer that the retrieved knowledge paths $P$ may contain incorrect or irrelevant information. To prevent such noisy knowledge from interfering with the LLM’s reasoning, we refrain from supplying P to the model. Instead, we employ chain-of-thought prompting to guide the LLM in leveraging its internal knowledge to generate answers. Example prompt can be seen in Appendix A.1.
4 Experiments
4.1 Experimental Settings
Datasets and Evaluation Metrics. To evaluate the performance of our proposed paradigm, We selected two open-source knowledge graphs and one self-constructed graph as external knowledge bases for experimentation. The open-source knowledge graphs are MOOC Q&A Yu et al. (2020) and Freebase Bollacker et al. (2008). The graph that we constructed ourselves is not yet publicly available and is referred to in this paper as the agricultural knowledge graph.
Freebase is a large-scale, semi-structured database supported by Google, designed to collect and connect information about millions of entities and their relationships worldwide. It covers a wide range of topics such as people, places, books, movies, music, etc., with each entity having a unique identifier. We conduct experiments on two QA sets with Freebase as external knowledge base: WebQSP Yih et al. (2016) and WebQuestions Berant et al. (2013).
For questions with multiple correct answers, we evaluate the results using two metrics: partial match and complete match. Partial match requires the experimental method to provide one or more correct answers, while complete match demands all correct answers.
Baselines. We compared our approach with six baseline methods: standard prompting (IO prompt) Brown et al. (2020), Self-Consistency (SC) Wang et al. (2023a), chain of thought prompting (CoT prompt) Wei et al. (2022), ToG Sun et al. (2024), PoG Chen et al. (2024), StructGPT Jiang et al. (2023) and KB-BINDER Li et al. (2023a). IO prompt, SC and CoT prompt are knowledge-free methods, used to measure how many questions LLMs can answer solely based on their internal knowledge. ToG, PoG, StructGPT and KB-BINDER represent previous state-of-the-art approaches in knowledge base question answering, serving as baselines to evaluate our knowledge retrieval method’s effectiveness.
Table 1: Results for MOOC Q&A. We report the accuracy of each method in answering "yes" or "no" correctly on the True/False question set, as well as the proportion of responses containing partially correct answers (partial match) and fully correct answers (complete match) on the Query question set.